OpenStudy (anonymous):

What is the simplified form of sqrt(98n^19)

sam (.sam.):

\[ \sqrt{98n^{19}} \\ \\ =\sqrt{98}\sqrt{n^{19}} \\ \\ =\sqrt{49}\sqrt{2}\sqrt{n^{19}}\] What's \(\sqrt{49}\)?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7 twice

sam (.sam.):

Good, now you get \[7\sqrt{2}\sqrt{n^{19}}\]

OpenStudy (anonymous):

yes, thank you :)